The 5th term of an arithmetic progression is three times the first term.find the sum of the first eight terms of the progression given that a quarter of the fifth term is 9
Answer by mananth(16946)

The 5th term of an arithmetic progression is three times the first term.find the sum of the first eight terms of the progression given that a quarter of the fifth term is 9
let first term be a
and common difference be d
t5 = 3a
(1/4)t5 =9
t5=36
3a=36
a=12
t5=a+4d
36=12+4d
24=4d
6=d
a=12,d=6
Sn = (n/2) (2a+(n-1)d)
S8=(8/2)(2*12+(7*6))
=4*(14+13)
=(4*27)
=108